what is the slope of the line passing throught the given points?
(2,3),(5,9)

Answer by checkley71(8403)

(2,3) are the x1 & y1 point (5,9) are the x2 & y2 points
(y2-y1)/(x2-x1)= slope or (9-3)/(5-2)=6/3 or 2 is the slope of the line passing through these points.